# Learning Visual Hierarchies in Hyperbolic Space for Image Retrieval

### v3

@@ 5.2 Generalization Evaluation
We evaluate the generalization of our method through image retrieval on unseen datasets and hierarchies. Similar to Table 1-2, Table 3 presents part-based same-class and hierarchical retrieval performance on the out-of-domain LVIS dataset [14], which contains 1,203 fine-grained object classes, including many rare objects absent from the training HierOpenImages dataset. Table 4 evaluates image retrieval performance on VOC and COCO datasets following the zero-shot object detection in [30] where ground-truth boxes are used as region proposals, and then predict via top-5 majority vote. Results demonstrate that our visual hierarchical learning significantly enhances model generalization on unseen datasets and image hierarchies.
@@
CLIP Cos Sim. 2.37 2.32 2.04 1.84
CLIP-euc Euc Ang. 19.36 18.64 16.14 14.52
CLIP-hyp Hyp Ang. 22.33 21.40 18.50 16.42
@@
CLIP Cos Sim. 8.57 7.77 5.70 4.64
CLIP-euc Euc Ang. 20.24 20.03 19.30 18.68
CLIP-hyp Hyp Ang. 22.41 22.29 21.70 21.10
@@
Recall $\uparrow$ CLIP Cos Sim. 18.30 32.42 45.42
CLIP-euc Euc Ang. 46.89 64.40 75.71
CLIP-hyp Hyp Ang. 47.57 64.71 75.72
OT Distance $\downarrow$ CLIP Cos Sim. 7.61 8.88 9.43
CLIP-euc Euc Ang. 5.74 7.39 8.43
CLIP-hyp Hyp Ang. 5.74 7.38 8.40
@@ B.1 Hyperparameters and training details
We employ the AdamW optimizer with parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$ and a learning rate of $2\times 10^{-5}$. Training was conducted using 8 $\times$ A10G Nvidia GPUs. For each model, we used the largest batch size that fit in memory: CLIP ViT was trained with a total effective batch size of 320, and MoCo-v2 with a total effective batch size of 800. Each model was fine-tuned for a single epoch on HierOpenImages dataset, taking approximately 26 hours for CLIP ViT and 18 hours for MoCo-v2. The embeddings are projected to dimension 128 in the final layer. The hyperbolic model has a learnable curvature parameter.
@@ B.2 Part-based Image Retrieval
HierOpenImages is built from the OpenImages dataset [24], which originally contains approximately 1.9 million images, 14 million bounding boxes and 600 labels. We create image-to-bounding box pairs, including one cross-image bounding box sample for each bounding box class in the image, and bounding box to bounding box pairs where at least $80\%$ of the smaller bounding box is contained in the larger bounding box.
@@
Figure 9: Example of our image retrieval interface built with Gradio [1]. The interface supports image selection, upload and retrieves results ranked by user-defined modes. Its modular design allows for easy integration of additional functionalities.
@@
We built our image retrieval interface using Gradio [1], as shown in Fig. 9. Input images can be selected from a linked image folder, where thumbnail images are displayed in an image gallery, or they can be directly uploaded by users. The retrieval results can be sorted by the hyperbolic angle relative to the input image or filtered using a user-defined threshold value, after which the results are ordered by their embedding norms. Additional functionalities can be easily integrated into the current pipeline.
@@ B.3 Generalization Evaluation
We evaluate the generalization capability of our model on the out-of-domain LVIS dataset [14], which is designed for large-scale ($\sim$1.2M bounding boxes) long-tail instance classification and segmentation. It has a highly imbalanced distribution of 1,203 object categories and contains 897 object categories that are absent from OpenImages [24]. Here are some examples of unseen hierarchies in LVIS [14] but not in OpenImges [24]: {table $\rightarrow$ tablecloth $\rightarrow$ ashtray $\rightarrow$ cigarette}, {backpack $\rightarrow$ strap $\rightarrow$ belt buckle}, {toy $\rightarrow$ teddy bear $\rightarrow$ thread $\rightarrow$ bobbin}, {sofa $\rightarrow$ blanket $\rightarrow$ quilt $\rightarrow$ bedspread}. The full list of these categories can be found in the appendix. Only display the first class of synonyms.
We construct the hierarchical evaluation set from the validation set of the LVIS dataset [14], following the similar process as constructing HierOpenImages. To construct the ground truth hierarchy tree, we use the pipeline as described in Sec. A. We empirically choose parameter $\textit{frequency}=5$ and $\textit{proportion}=5\%$ in our experiments. This adjustment is necessary because the LVIS dataset [14] has very unbalanced classes.
@@
CLIP-hyp 73.37 72.59 69.84 68.62
CLIP-euc 67.83 68.37 67.12 66.04
CLIP-hyp 70.47 70.29 67.89 66.38
CLIP-euc 64.98 64.49 63.51 63.10
CLIP - 26.73 25.95 23.69 22.70
@@
CLIP-hyp 66.02 66.63 66.50 65.91
CLIP-euc 65.38 65.70 66.01 65.79
CLIP-hyp 63.00 63.78 63.65 63.44
CLIP-euc 60.33 60.73 61.43 61.38
CLIP - 47.52 46.60 43.50 42.08
Table 5: Part-based same-class image retrieval evaluation. Cross-image ✓indicates models fine-tuned on entailment relationships both within and across images at the category level, while ✗ represents models fine-tuned without cross-image sampling. The evaluation setup is the same as Table 1 in the main paper. The best results are marked in bold, and the second-best results are in blue.
@@
CLIP-hyp 74.06 82.34 88.79
CLIP-euc 72.96 80.98 87.97
CLIP-hyp 73.52 81.58 88.10
CLIP-euc 71.34 78.89 86.04
Recall $\%\uparrow$ CLIP - 51.33 64.28 77.02
CLIP-hyp 16.36 19.27 22.21
CLIP-euc 17.00 20.00 22.45
CLIP-hyp 16.95 19.77 22.55
CLIP-euc 18.43 21.25 23.53
OT Distance $\downarrow$ CLIP - 23.81 24.60 25.03
Table 6: Part-based hierarchical evaluation of parent-to-child image retrieval. The OT distance is defined in Sec. 3.3 (main paper), and the evaluation setup follows Table 2 (main paper). The best results are marked in bold, and the second-best results are in blue.
@@ Appendix C Ablation Studies
In this ablation study, we evaluate the impact of cross-image scene-to-object samples. We fine-tuned the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), excluding any cross-image image-to-bounding-box samples. Fig. 4 in the main paper shows that the hyperbolic model CLIP-hyp fine-tuned without cross-image samples significantly outperforms the Euclidean model CLIP-euc and even surpasses the model trained with additional cross-image samples. This indicates that training in hyperbolic space enhances the model’s ability to recognize visually dissimilar entailment pairs.
In this supplementary material, we further evaluate these models on part-based same-class and hierarchical image retrieval tasks, as shown in Table. 5-6. The best results are highlighted in bold and the second-best results are shown in blue. The results clearly indicate that cross-image sampling improves image retrieval performance. Notably, the hierarchical retrieval results align with Fig. 4 in the main paper. Specifically, the hyperbolic model trained without cross-image samples outperforms the Euclidean model trained with cross-image samples in 5 out of 6 cases, as shown in Table 6.

### v4

@@ 5.2 Generalization Evaluation
We evaluate the generalization of our method through image retrieval on unseen datasets and hierarchies. Similar to Table 1-2, Table 3 presents part-based same-class and hierarchical retrieval performance on the out-of-domain LVIS dataset [gupta2019lvis], which contains 1,203 fine-grained object classes, including many rare objects absent from the training HierOpenImages dataset. Table 4 evaluates image retrieval performance on VOC and COCO datasets following the zero-shot object detection in [pal2024compositional] where ground-truth boxes are used as region proposals, and then predict via top-5 majority vote. Results demonstrate that our visual hierarchical learning significantly enhances model generalization on unseen datasets and image hierarchies.
@@
CLIP Cos Sim. 15.00 14.22 11.91 10.48
CLIP-euc Euc Ang..* 28.45 26.76 22.65 20.43
CLIP-hyp Hyp Ang..* 28.84 27.02 22.87 20.53
@@
CLIP Cos Sim. 28.37 27.35 24.99 23.36
CLIP-euc Euc Ang..* 28.84 27.02 22.65 20.43
CLIP-hyp Hyp Ang..* 30.27 29.53 27.83 26.81
@@
Recall $\uparrow$ CLIP Cos Sim. 42.40 57.22 67.98
CLIP-euc Euc Ang..* 53.54 68.77 78.61
CLIP-hyp Hyp Ang..* 53.87 69.13 78.85
OT Distance $\downarrow$ CLIP Cos Sim. 6.25 7.89 8.76
CLIP-euc Euc Ang..* 5.49 7.04 8.05
CLIP-hyp Hyp Ang..* 5.56 7.13 8.13
@@ B.1 Hyperparameters and training details
We employ the AdamW optimizer with parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$ and a learning rate of $2\times 10^{-5}$. Training was conducted using 8 $\times$ A10G Nvidia GPUs. For each model, we used the largest batch size that fit in memory: CLIP ViT was trained with a total effective batch size of 640, and MoCo-v2 with a total effective batch size of 1984. Each model was fine-tuned on HierOpenImages dataset. The embeddings are projected to dimension 128 in the final layer. The hyperbolic model has a learnable curvature parameter.
@@ B.2 Part-based Image Retrieval
HierOpenImages is built from the OpenImages dataset [kuznetsova2020open], which originally contains approximately 1.9 million images, 14 million bounding boxes and 600 labels. We create image-to-bounding box pairs, including one cross-image bounding box sample for each bounding box class in the image, and bounding box to bounding box pairs where at least $80\%$ of the smaller bounding box is contained in the larger bounding box.
@@
Gated Entailment Angle Metric.
In this work, all entailment-angle thresholds are measured in radians. We use a threshold of 2 for MoCo and 2.5 for CLIP in small-k same-class classification, and a threshold of 1 for both models in large-k hierarchical retrieval.
@@
Figure 9: Example of our image retrieval interface built with Gradio [abid2019gradio]. The interface supports image selection, upload and retrieves results ranked by user-defined modes. Its modular design allows for easy integration of additional functionalities.
Gated Entailment Angle Metric.
Vision Encoder Model Metrics Top-5 Top-10 Top-50 Top-100
Child-to-Parent
CLIP ViT CLIP-euc Cos Sim. 75.20 74.32 71.81 70.21
CLIP-hyp Cos Sim. 76.82 75.53 72.46 70.50
CLIP-euc Euc Ang. 75.35 74.23 72.45 70.78
CLIP-hyp Hyp Ang. 76.95 75.97 72.94 71.04
CLIP-euc Euc Ang..* 75.63 74.65 72.25 70.52
CLIP-hyp Hyp Ang..* 77.28 75.91 72.85 70.94
MoCo-v2 MoCo-euc Cos Sim. 67.93 67.67 66.20 65.04
MoCo-hyp Cos Sim. 69.00 68.37 66.26 64.96
MoCo-euc Euc Ang. 47.00 46.64 43.36 41.00
MoCo-hyp Hyp Ang. 49.91 48.93 43.73 40.95
[…]
Parent-to-Child
CLIP ViT CLIP-euc Cos Sim. 73.73 72.52 69.97 68.70
CLIP-hyp Cos Sim. 74.38 73.20 70.67 69.28
CLIP-euc Euc Ang. 72.59 71.58 69.49 68.46
CLIP-hyp Hyp Ang. 73.80 72.89 70.67 69.52
CLIP-euc Euc Ang..* 73.74 72.57 70.04 68.78
CLIP-hyp Hyp Ang..* 74.48 73.27 70.78 69.46
MoCo-v2 MoCo-euc Cos Sim. 66.68 65.90 64.20 63.25
MoCo-hyp Cos Sim. 66.80 65.88 64.07 63.11
MoCo-euc Euc Ang. 67.21 66.23 63.47 61.85
MoCo-hyp Hyp Ang. 67.53 66.14 62.50 60.82
MoCo-euc Euc Ang..* 68.87 67.94 65.64 64.42
MoCo-hyp Hyp Ang..* 69.50 68.14 65.67 64.46
Table 5: Part-based same-class image retrieval evaluation using different metrics. Cosine similarity (Cos Sim.) is the pretrained metric, entailment angle (Euc or Hyp Ang.) is the finetuned metric, and * represents a gated entailment angle metric which combines the strengths of both.
In Table 5, we report retrieval performance using three metrics for completeness: cosine similarity (the model’s pretrained metric), entailment angle (the finetuned metric), and a gated entailment angle metric (Ang.*), which combines the strengths of both.
The gated approach first uses the entailment angle as a cone-shaped filter to retain only candidates that satisfy the desired hierarchical direction (e.g., retrieving a parent or child image) and discards those that violate the entailment relationship. Cosine similarity is then applied within this gated set to provide fine-grained ranking. This two-stage procedure ensures that retrieval respects semantic hierarchy while leveraging the pretrained model’s strong embedding geometry.
A key finding is that fine-tuning with the entailment-angle loss not only improves performance under the entailment angle metric itself but also enhances the model’s original pretrained cosine-similarity performance. This suggests that the fine-tuning introduces meaningful hierarchical structure into the representation space in a way that strengthens the model’s pretrained retrieval capability. This observation aligns with the improvements observed in the zero-shot retrieval or classification results on LVIS, VOC, and COCO in the main paper.
Our intuition is that for models with strong pretrained priors or similar training setups (e.g., MoCo-v2, which is trained on image-only data with invariance to random crops), entailment angle fine-tuning primarily reshapes the coarse organization of the embedding space, aligning representations with the angle-based entailment structure. However, because these models inherit strong inductive priors from massive pretraining, small finetuning cannot fully reorient the fine-grained similarity relationships. The gated metric effectively leverages these complementary strengths. The entailment angle enforces the correct hierarchical region, while cosine similarity provides accurate ranking within that region. By combining directional gating with pretrained metric ordering, the hybrid approach balances the new hierarchical signal with the stability of the pretrained embedding, yielding a more robust and reliable retrieval strategy.
@@
We built our image retrieval interface using Gradio [abid2019gradio], as shown in Fig. 9. Input images can be selected from a linked image folder, where thumbnail images are displayed in an image gallery, or they can be directly uploaded by users. The retrieval results can be sorted by the hyperbolic angle relative to the input image or filtered using a user-defined threshold value, after which the results are ordered by their embedding norms. Additional functionalities can be easily integrated into the current pipeline.
@@ B.3 Generalization Evaluation
We evaluate the generalization capability of our model on the out-of-domain LVIS dataset [gupta2019lvis], which is designed for large-scale ($\sim$1.2M bounding boxes) long-tail instance classification and segmentation. It has a highly imbalanced distribution of 1,203 object categories and contains 897 object categories that are absent from OpenImages [kuznetsova2020open]. Here are some examples of unseen hierarchies in LVIS [gupta2019lvis] but not in OpenImges [kuznetsova2020open]: {table $\rightarrow$ tablecloth $\rightarrow$ ashtray $\rightarrow$ cigarette}, {backpack $\rightarrow$ strap $\rightarrow$ belt buckle}, {toy $\rightarrow$ teddy bear $\rightarrow$ thread $\rightarrow$ bobbin}, {sofa $\rightarrow$ blanket $\rightarrow$ quilt $\rightarrow$ bedspread}. The full list of these categories can be found in the appendix. Only display the first class of synonyms.
We construct the hierarchical evaluation set from the validation set of the LVIS dataset [gupta2019lvis], following the similar process as constructing HierOpenImages. To construct the ground truth hierarchy tree, we use the pipeline as described in Sec. A. We empirically choose parameter $\textit{frequency}=5$ and $\textit{proportion}=5\%$ in our experiments. This adjustment is necessary because the LVIS dataset [gupta2019lvis] has very unbalanced classes.
@@
CLIP-hyp 77.28 75.91 72.85 70.94
CLIP-euc 75.63 74.65 72.25 70.52
CLIP-hyp 73.56 71.57 68.25 66.60
CLIP-euc 73.32 71.26 67.33 65.67
CLIP - 53.04 51.69 47.87 45.79
@@
CLIP-hyp 74.48 73.27 70.78 69.46
CLIP-euc 73.74 72.57 70.04 68.78
CLIP-hyp 71.87 69.58 65.31 63.95
CLIP-euc 71.61 69.00 64.69 63.28
CLIP - 69.94 67.74 62.50 60.21
Table 6: Part-based same-class image retrieval evaluation. Cross-image ✓indicates models fine-tuned on entailment relationships both within and across images at the category level, while ✗ represents models fine-tuned without cross-image sampling. The evaluation setup is the same as Table 1 in the main paper. The best results are marked in bold.
@@
CLIP-hyp 77.00 85.75 91.89
CLIP-euc 76.46 85.08 91.38
CLIP-hyp 75.66 84.51 91.13
CLIP-euc 75.22 84.15 91.10
Recall $\%\uparrow$ CLIP - 66.63 77.15 86.57
CLIP-hyp 14.96 17.57 20.76
CLIP-euc 15.65 18.15 21.09
CLIP-hyp 15.98 18.27 21.03
CLIP-euc 16.19 18.46 21.02
OT Distance $\downarrow$ CLIP - 21.31 22.74 23.79
Table 7: Part-based hierarchical evaluation of parent-to-child image retrieval. The OT distance is defined in Sec. 3.3 (main paper), and the evaluation setup follows Table 2 (main paper). The best results are marked in bold.
@@ Appendix C Ablation Studies
In this ablation study, we evaluate the impact of adding cross-image scene-to-object samples. We fine-tuned the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), excluding any cross-image image-to-bounding-box samples. The detailed results for same-class retrieval and hierarchical retrieval are presented in Table 67, with the best scores highlighted in bold. Overall, the results clearly show that incorporating cross-image samples further improves image retrieval performance.